OpenStudy (anonymous):

If the midpoint between (x, 3) and (9, 9) is (7, 6), find the value of x.?

OpenStudy (zpupster):

x+9/2=7

OpenStudy (anonymous):

It said midpoint of those two points is (7,6) x + 9/2 = 7 (x2 - x1/2, y2 - y1/2) is mid point formula x + 9 = 14 hope u get from there

OpenStudy (anonymous):

i dont get it at all

OpenStudy (anonymous):

sorry I worte the formula wrong

OpenStudy (anonymous):

Midpoint of two points is given by ((x2 + x1)/2 , (y2 + y1)/2)

OpenStudy (anonymous):

@zpupster 7 is not an answer

OpenStudy (anonymous):

so here x1 = x and x2 = 9 given midpoint x - cordinate is 7 now according to the formula (9 + x)/2 = 7

OpenStudy (zpupster):

ok you have to solve for x first mult both sides by 2 x+9/2*2=7*2 x + 9 = 14 then subtract both sides by 9 x+9-9 =14-9 x=5

OpenStudy (anonymous):

9 +x = 14 x = 14 - 9 x = 5

OpenStudy (anonymous):

thank u

OpenStudy (zpupster):

i did not say 7 is the answer i plugged the numbers in the midpoint formula